The perimeter of a rectangular field is 504 yards. It's length is 6 yards shorter than twice it's width? What is the area?
Answer by algebrahouse.com(1659)

"The perimeter of a rectangular field is 504 yards. It's length is 6 yards shorter than twice it's width? What is the area?"

x = width
2x - 6 = length {six yards shorter than twice the width}

2(x) + 2(2x - 6) = 504 {two times width + two times length = perimeter}
2x + 4x - 12 = 504 {used distributive property}
6x - 12 = 504 {combined like terms}
6x = 516 {added 12 to both sides}
x = 86 {divided both sides by 6}
2x - 6 = 166 {substituted 86 for x into 2x - 6}

width = 86 yards
length = 166 yards
area = 14,276 yds^2 {area is length x width}
www.algebrahouse.com